### v2

# Multimodal LLM Guided Exploration and Active Mapping using Fisher Information

Abstract
We present an active mapping system that could plan for long-horizon exploration goals and short-term actions with a 3D Gaussian Splatting (3DGS) representation. Existing methods either did not take advantage of recent developments in multimodal Large Language Models (LLM) or did not consider challenges in localization uncertainty, which is critical in embodied agents. We propose employing multimodal LLMs for long-horizon planning in conjunction with detailed motion planning using our information-based algorithm. By leveraging high-quality view synthesis from our 3DGS representation, our method employs a multimodal LLM as a zero-shot planner for long-horizon exploration goals from the semantic perspective. We also introduce an uncertainty-aware path proposal and selection algorithm that balances the dual objectives of maximizing the information gain for the environment while minimizing the cost of localization errors. Experiments conducted on the Gibson and Habitat-Matterport 3D datasets demonstrate state-of-the-art results of the proposed method.
1 Introduction
Being able to autonomously explore and map an environment while localizing within that map is a core skill for a mobile robot. This ability could empower embodied artificial intelligence systems with effective 3D scene understanding through use in conjunction with open-vocabulary semantic segmentation or vision-language features [91, 68, 55, 87] and can be used as a basis for language-specified robotics tasks [25, 63, 37, 26]. This task is challenging because it requires general knowledge of the typical layout of an environment to identify coarse targets for long-term exploration and the ability to accurately plan each step to ensure localization accuracy while maximizing information gain or ‘surprise’ when traversing the environment.
In this paper, We propose to use 3D Gaussians as our scene representation and tackle this problem in two phases: long-horizon planning with an LLM and detailed motion planning with our information-based approach by quantifying the Fisher Information for 3DGS parameters and current localization states. We prompt a multimodal LLM with our current map state, trajectory, and candidates for long-term exploration frontiers to exploit the LLM’s world knowledge about scene layouts. We then follow the suggestion from the LLM by choosing paths that maximize the Expected Information Gain calculated using the Fisher Information about poses and 3D Gaussians.
Figure 1: An overview about the interaction with a multimodal LLM. The image on the left is a bird’s-eye view rendering from our 3DGS representation. A detailed example of the dialog can be found in the supplementary.
Classical active mapping methods often define objectives to reduce the uncertainty of the state estimate based on its covariance [20, 58, 5, 78, 11, 1]. The covariance is readily available for classical systems which use filters to update the state estimate. However, many recent mapping systems use non-linear optimization to update the state estimate, making the covariance difficult to obtain. In particular, systems [44, 83, 32, 27] with 3DGS [33] for the scene representation have been developed for high-fidelity rendering of novel views of the scene. Previous methods attempted to quantify the uncertainty of radiance fields for reconstructing scenes from given data [66, 67, 72, 23] for active view selection [51, 31, 72, 36] and active reconstruction [84] or mapping [85] of scenes with given localization, and for 3D reconstruction of small scenes with an inward-facing camera [88]. However, all the prior methods only model the uncertainty of the scene representation, ignoring another primary source of error for a practical mapping system: the risk of localization failures. Our proposed system could address not only Information Gain during exploration but also consider localization accuracy when the robot explores unknown and texture-less regions.
In terms of long-horizon planning for robot exploration, classical approaches such as frontier-based exploration and A* algorithms are still used in active mapping systems for their efficiency and simplicity [70, 4, 77, 71, 34]. However, algorithms that use only simple heuristics cannot determine the information gain, nor estimate 3D geometry, limiting their objective to simply improving coverage or having minimal travel distances. To address this issue, learnable approaches have been trained for exploration tasks [22, 59, 38, 46], but these methods are trained for a limited scene distribution and are not natively designed for a 3D Gaussian representation. Recently, LLMs and Vision Language Models (VLM) have demonstrated extraordinary abilities in visual grounding and logical reasoning and have been studied with various robotic tasks [62]. However, these works either focus on other tasks [9, 16, 14], especially tasks with language inputs [26, 74] and navigation [8] or did not take advantage of photo-realistic 3D representation for multimodal LLMs. On the other hand, Neural Radiance Field and 3D Gaussian Splatting have shown the exciting ability to reconstruct accurate 3D scenes just using posed images. Later works [44, 32] enabled online training with localizations of images. However, the 3DGS model reconstructs the scene from a pure optimization perspective without prior knowledge of the scene and only passively works on RGB-D video inputs. In contrast, we can actively select the best paths for exploration to improve the quality of both reconstruction and localization.
Figure 2: Scatter Plot of EIG vs. PSNR We plot the EIG and PSNR at sampled poses in the Cantwell scene of the Gibson dataset. The figure corroborates the intuition that the robot expects to gain little information (low EIG) at well reconstructed region (high PSNR) and gain much information (high EIG) at a poorly reconstructed region (low PSNR).
We argue that we can leverage the best of both worlds. We can use high-quality bird’s-eye view rendering from a 3D Gaussian Splatting model as the input of the multimodal LLM to help identify long-horizon exploration goals from frontiers in our scene representation. In this way, we could seamlessly integrate multimodal LLM into a path-planning model to provide knowledge of the typical layout of a scene and potential regions of interest for exploration. Besides, we also alleviate the issue by asking the multimodal LLM to provide precise coordinates, which require precise 3D understanding and are dependent on the scale and relative location of the scene. Given the long-horizon exploration goal from the output of the multimodal LLM, we formulate our path planning problem as an active learning problem and propose multiple paths leading to the exploration goal. We consider the expected information gain (EIG) along the planned trajectory with Fisher Information. As shown in Fig. 2, our EIG metric could reflect the ‘surprise’ and correlate well with the rendering quality on candidate views without actually taking a picture at the candidate location. We also derived localization uncertainty on candidate paths with Cramér–Rao bound to balance the dual problem of exploration and localization. Notably, we compute the full Fisher Information matrix for localization parameters without any approximation with our efficient CUDA implementation. Thus, finding the path that most improves scene reconstruction quality when taking observations during exploration while ensuring the agent still has accurate localization. To validate our approach, we evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D [56] datasets quantitatively and qualitatively. We show superior reconstruction quality in various metrics compared to several baselines and recent state-of-the-art methods [7, 53, 22, 85, 82]. Our contributions can be summarized as follows:
  • We present an active mapping system that could autonomously explore the environment and extensively compare our system with previous approaches. To the best of our knowledge, we are the first active mapping system with 3D Gaussian representation that is not dependent on ground truth camera pose readings.
  • We provide a way to leverage the zero-shot long-horizon planning ability of LLM into our active mapping system seamlessly.
  • We introduce localization uncertainty with active mapping systems and effectively balance the information gain for exploration and the cost of possible localization errors.
2 Related work
Active Mapping and Localization
Efficiently exploring an environment in order to map it while being able to localize in that map is a fundamental problem in robotics. There are many methods that address aspects of this problem separately – Simultaneous Localization and Mapping (SLAM) methods [48, 18, 47, 3, 30, 49, 17, 75] address the mapping and localization. Many exploration methods [22, 85, 10] address the exploration and mapping aspects while assuming poses are provided. Active SLAM methods [70, 4, 77, 71] consider both of these problems. The exploration in these systems is usually driven by a measure of uncertainty [39, 54]; specific utility functions are often drawn from either Information Theory (IT) [65], or the Theory of Optimal Experimental Design (TOED) [52]. Recently learning-based approaches have been developed for active mapping and localization, Active Neural SLAM [7] learns policies to drive exploration and estimate the agent pose, more similarly to us NARUTO [21] uses a SLAM backbone to estimate the pose and uses uncertainty to drive exploration. However, they only consider the reconstruction uncertainty, whereas we choose long-horizon planning goals and consider the localization uncertainty as well as the reconstruction uncertainty when planning the best path to the long-horizon goal.
Uncertainty quantification for radiance fields
The vast majority of previous work on uncertainty quantification for radiance fields has been for post-processing scenes [66, 67, 72, 23], view selection [51, 31, 72, 36] or active view selection [51, 72, 31], all of which assume the input images are posed. Active neural mapping [85] uses neural variability, that is the prediction robustness against random weight perturbations, as an estimate of uncertainty to actively map a scene with ground truth poses provided. Fisher-RF [31] also performs active scene mapping with ground truth poses provided, based on an approximation of the Fisher Information of views along candidate paths. Zhan et al. [88] perform active reconstruction without ground truth camera poses. However, they only evaluate small-scale scenes and limit the camera trajectories to be inwards facing and only model scene uncertainty, not localization uncertainty, which is a key consideration for active mapping and exploration.
Robot Planning with Foundation Models
LLMs and VLMs have been used widely in robotics as for language-specified tasks, either as high-level planners in conjunction with other methods for low-level control [28, 26, 40, 69, 2, 57, 16, 15], to create intermediate representations which can be planned over [29] or directly outputting actions [90, 41]. Outside of language-driven tasks, VLMs and LLMs have been used in robotics for providing rewards to drive exploration in Reinforcement Learning [76], and for visual localization [45]. More similar to our task, LLMs and VLMs have also been used for goal-driven navigation, such as searching for specific objects [86, 60, 13, 64].
3 Method
Figure 3: An Illustration of Our Active Mapping System Our method first identifies long-horizon targets with a multimodal LLM by using novel-view synthesis from our 3DGS representation. Our information-based path proposal method then proposes and selects short-term action sequences based on the Fisher Information about the 3D Gaussian parameters. The best path and action along the path is selected with respect to both the information gain and localization accuracy.
We divide the scene exploration task into two phases: (a) long-horizon planning for the coarse direction of exploration that leads to better coverage and understanding (Sec. 3.2), and (b) detailed trajectory proposal (Sec. 3.3) and path selection (Sec. 3.4) that aims to improve 3D reconstruction and localization from a geometrical perspective. The former task relies more on semantic information and prior knowledge about the possible layout of the environment, which is suitable for a generalist model with prior distributions on the scene. The latter task, however, involves detailed motion planning and is better addressed by considering information gain on fine-level geometries. Fig. 3 shows an overview of our method.
3.1 Preliminary
In 3D Gaussian Splatting (3DGS) [33], the scene is represented by a set of 3D Gaussians whose color and opacity are learned via a rendering loss. A rendered image can be produced by projecting the Gaussians to 2D and using $\alpha$-blending for the $N$ ordered points on the 2D splat that overlaps each pixel. The Jacobian of the current camera pose ${\bf x}$ with respect to the parameters of 3D Gaussians can be computed by defining the (left) partial derivative on the manifold [44]:
$\frac{\mathcal{D}{f({\bf x})}}{\mathcal{D}{{\bf x}}}\triangleq\lim_{\tau\to 0}\frac{\text{Log}(f(\text{Exp}(\tau)\circ{\bf x})\circ f({\bf x})^{-1})}{\tau}~{},$ (1)
where $\tau\in\mathfrak{se}(3)$, $\circ$ is a group composition, and Exp and Log are the exponential and logarithmic mappings between Lie algebra and Lie Group.
Fisher Information is a measurement of the information that a random variable ${\bf y}$ carries about an unknown parameter ${\bf w}$ of a distribution that models ${\bf y}$. In the problem of novel view synthesis, we are interested in measuring the observed information of a radiance field with parameters ${\bf w}$ at a camera pose ${\bf x}$ using the negative log-likelihood of the image observation ${\bf y}$ taken from that pose:
$-\log p({\bf y}|{\bf x};{\bf w})=({\bf y}-f({\bf x},{\bf w}))^{T}({\bf y}-f({\bf x},{\bf w})),$ (2)
where $f({\bf x},{\bf w})$ is the rendering model. Under regularity conditions [61], the Fisher Information of $-\log p({\bf y}|{\bf x};{\bf w})$ is the Hessian of Eq. 2 with respect to ${\bf w}$, denoted ${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]$. In our formulation ${\bf w}$ is a tensor for all the 3D Gaussian Parameters $R^{N\times 14}$, where $N$ is the number of Gaussians, and ${\bf y}\in R^{\it{h}\times\it{w}\times\it{c}}$ is the RGB-D observation.
FisherRF [31] addressed the active view selection problem that starts with a training set of views $D^{train}$ and aims to select the next best view from a set of candidate $SE(3)$ camera poses ${\bf x}^{acq}_{i}~{}\in~{}D^{pool}$ without obtaining the image ${\bf y}^{acq}_{i}$ at the camera pose ${\bf x}^{acq}_{i}$. The next best view is chosen by finding:
$\operatorname*{arg\,max}_{{\bf x}^{acq}_{i}\in D^{pool}}\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]\;{\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}\right),$ (3)
where ${\bf w}$ is the initial estimate of model parameters using the current training set $D^{train}$. ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]^{-1}$ can be computed by summing the Hessians of model parameters across all different views in $\{D_{train}\}$ before inverting. The key of this algorithm is that the Fisher Information ${\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]$ does not depend on the label ${\bf y}^{acq}_{i}$ of the acquisition sample ${\bf x}^{acq}_{i}$. Therefore, it is feasible to compute the Expected Information Gain (EIG) before visiting the potential view candidate ${\bf x}^{acq}_{i}$. However, the number of optimizable parameters is typically more than 20 million, which means it is impossible to compute without sparsification or approximation. In practice, FisherRF [31] applies a Laplace approximation [12, 42] that approximates the Hessian matrix with its diagonal values plus a log-prior regularizer $\lambda I$
${\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]\simeq\operatorname{diag}(\nabla_{\bf w}f({\bf x},{\bf w})^{T}\nabla_{\bf w}f({\bf x},{\bf w}))+\lambda I.$ (4)
3.2 Long Horizon Exploration with Foundation Model
We propose using a multimodal LLM to perform the task of long-horizon planning as a zero-shot planner for the possible paths and leave the detailed path planning for a closed-form uncertainty-aware motion planning algorithm that computes 3D geometries.
As we consider an agent moving in a 2D action plane (e.g., a ground robot), we leverage the view synthesis ability of 3DGS to create expressive bird’s-eye view renderings that could provide an overview of the environment. An occupancy grid on the motion plane of our robot is created using our 3D Gaussian representation, which can be used to identify the frontiers of the current environment. The frontiers are defined as points on the boundary between free space and unobserved space (not marked as free or occupied) in our occupancy map. Each frontier can be defined as a set of neighboring points on the action plane $Q=\{\phi({\bf x}_{i})...\}$, where $\phi(.)$ is the projection from the world frame to the bird’s-eye view. The agent has multiple choices on the frontiers, especially at the early stages of exploration. We apply Chain-of-Thought Prompting [80] to encourage the multimodal LLM to provide analysis on the candidate frontiers first before selection. We provide contextual information such as the description of the task, the total steps allowed for exploration, and our current step in the textual prompt. The multimodal LLM is also allowed to decide whether the robot should not go to a frontier and instead focus on improving existing regions. We embed possible choices of frontiers along with the current location and previously visited trajectory in the bird’s-eye view rendering of the environment along with a text prompt describing the task and contextual information as the input for the multimodal LLM. In this way, we avoid possible infeasible destinations for motion planning. For example, the unexplored regions we have no reconstruction for at this point might be unreachable from the current space. We then parse the long-horizon exploration goal from the reply of LLM. We use GPT-4o [50] for our long-horizon planning task, but our method is agnostic to the underlying multimodal LLM as the inputs for the multimodal LLM are text prompts and a bird’s-eye view rendering. Fig. 1 provides an overview of our interaction with the multimodal LLM. Detailed examples can be found in the Supplementary.
3.3 Information-driven Path Proposal
After identifying a long horizon with the multimodal LLM, we form an initial set of candidate poses $\mathcal{T}_{I}$ by sampling points in the coarse region of interest from the LLM. If there are no unvisited boundaries or the multimodal LLM suggests we do not need to explore the frontiers, we sample poses across the free space to form $\mathcal{T}_{I}$. We then evaluate the Expected Information Gain (EIG) for each pose ${\bf x}^{acq}_{i}\in\mathcal{T}_{I}$, given by
$\mathtt{EIG}({\bf x}^{acq}_{i})=\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}^{acq}_{i}|{\bf x}^{acq}_{i},{\bf w}]\;\mathcal{I}({\bf w})^{-1}\right),$ (5)
as a preliminary selection metric to form our final candidate target poses set $\mathcal{T}_{F}$. Thus, we can identify multiple coarse directions for explorations where we can propose multiple paths for detailed path planning and path selection. FisherRF [31] uses ${\mathbf{H}}^{\prime\prime}[{\bf w}|D^{train}]$ as an approximation for the observed Fisher Information $\mathcal{I}({\bf w})$ by computing the Hessians on the training set. This is also known as empirical Fisher Information, whose limitations have been widely discussed by Kunstner et al. [35] and Marten et al. [43]. In most scenarios, this is a reluctant design choice because the distribution of ${\bf x}\sim p({\bf x})$ is unknown (i.e., the distribution of all possible images). However, ${\bf x}\sim p({\bf x})$ in our case is tractable because it represents the possible locations where we can take an observation for the environment, a.k.a. the free space of our map. Therefore, unlike FisherRF [31] as described in Eq. 3, we propose to use Monte-Carlo sampling to compute the Fisher Information of the current model
$\mathcal{I}({\bf w})=\mathbb{E}_{{\bf x}\sim p({\bf x})}\left[{\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]\right]\simeq\sum_{k=1}^{N}{\mathbf{H}}^{\prime\prime}[{\bf y}_{k}|{\bf x}_{k},{\bf w}],$ (6)
where ${\bf x}_{k}$ is drawn from a uniform distribution of camera poses in the free space of the current map which we use to approximate $p$. Besides, we also uniformly initialize 3D Gaussians in the space, which will be subsequently updated with rendering losses for visited regions. In Fig. 2, we show the relationship between PSNR and EIG on sampled poses in the Cantwell scene. The result aligns with our expectation that EIG should decrease as PSNR increases. We also show some cases on the lower-left and upper-right of the scatter plot. For the lower left capture, even though the left part is poorly reconstructed, most of the view is occupied by a textureless wall, leading to a low EIG. For the upper right capture, even though the scene has a moderate reconstruction, the content is rich, so our algorithm returns a high EIG.
Finally, we compute paths towards each pose in $\mathcal{T}_{F}$ with the A* algorithm [24] using the occupancy map, selecting which path to follow as described in Section 3.4. The path can be defined as an ordered set of camera poses from the current location $x_{t}$ at exploration step $t$ to the frontier points $x_{T}^{j}$ from the long horizon planning.
$\mathcal{P}_{j}=\{x_{t+1}^{j},\dots,x_{T}^{j}\}$ (7)
ANS [7] Active-INR [85] UPEN [22] ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative Comparison for Final Scene Reconstruction on Gibson Dataset Greigsville (top) and Ribera (bottom) scenes. We provide top-down rendering for different methods. Note that UPEN and Active-INR use GT pose in this visualization.
UPEN [22] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Qualitative Comparison for Final Scene Reconstruction on Habitat-Matterport 3D Dataset mscxX4KEBcB (top), oPj9qMxrDEa (middle) and QKGMrurUVbk (bottom) scenes. We provide top-down rendering for different methods.
3.4 Short-term Path Selection with Localization Uncertainty
The EIG for 3D Gaussian parameters along a path $\mathcal{P}_{j}$ can be computed as the sum over the path of the following term [31]:
$\mathtt{EIG}_{\mathcal{P}_{j},i}({\bf x}_{i})=\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}_{i}|{\bf x}_{i},{\bf w}]\;\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})^{-1}\right)$ (8)
where $\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})$ takes the mutual information along the path into account as follows
$\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})={\mathbf{H}}^{\prime\prime}[{\bf w}]+\sum_{{\bf x}_{t}\in\mathcal{P}_{j},t<i}{\mathbf{H}}^{\prime\prime}[{\bf w}|{\bf x}_{t}].$ (9)
[…]
where ${\mathbf{H}}^{\prime\prime}[{\bf x}]$ is short for $-\nabla^{2}\log p(w)$ for clarity. If solely maximizing the $\mathtt{EIG}$, the robot will be more likely to explore unvisited regions. However, exploring regions that have not been well reconstructed also means the agent would have the risk of worse localization accuracy due to noise and ambiguities in the unreconstructed regions during pose optimization. The cost of localization must thus be considered during path planning to balance the importance of exploring new environments with maintaining localization accuracy. We propose to use Fisher Information as a measurement for the localization uncertainty that is also necessary for effective path planning for active Mapping algorithms. During optimization, we essentially optimize on the logarithmic mapping of $\tau_{i}\triangleq\text{Log}({\bf x}_{i})$ of our camera pose. By the Cramér–Rao bound, the covariance of $\tau_{i}\in\mathfrak{se}(3)$ can be lower-bounded with the inverse of Fisher Information matrix $\mathcal{I}(\tau_{i})$:
$\mathrm{Cov}(T(\hat{\tau_{i}}))\geq\mathcal{I}(\tau_{i})^{-1}$ (10)
where $T(\tau_{i})$ is an unbiased estimator for $\tau$ solved by iteratively optimizing photo-metric loss. Hence, we can define the localization cost $C_{\textit{local}}$ at a pose ${\bf x}_{i}$ in terms of $\tau_{i}$ as:
$C_{\textit{local}}(\tau_{i})=\log\det(\nabla_{\tau_{i}}f(\tau_{i},{\bf w})^{T}\nabla_{\tau_{i}}f(\tau_{i},{\bf w}))$ (11)
Matsuki et al. [44] computed the Jacobians of the camera pose with respect to the mean and covariances of each gaussian $\frac{\partial{\boldsymbol{\mu}_{I}}}{\partial{{\bf x}}}$ and $\frac{\partial{\boldsymbol{\Sigma}_{I}}}{\partial{{\bf x}}}$. However, we need to compute the Jacobian of $\tau_{i}$ with respect to the rendering output:
$\nabla_{\tau_{i}}f(\tau_{i},{\bf w})=\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\tau_{i}}}=\begin{bmatrix}\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\mu}_{I}}}&\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\Sigma}_{I}}}\end{bmatrix}\begin{bmatrix}\frac{\mathcal{D}{\boldsymbol{\mu}_{C}}}{\mathcal{D}{\tau_{i}}}\\ \frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{\tau_{i}}}\end{bmatrix}$ (12)
Unlike 3DGS parameters, our computation of the localization cost can be directly computed on the full Fisher Information matrix without using the Laplace approximation [12, 42]. That is because the full Jacobian $\nabla_{\tau_{i}}f(\tau_{i},{\bf w})$ is linear to the number of Gaussians and can be efficiently computed with our CUDA implementation.
Without loss of generality, the path of exploration can be selected by minimizing the total cost for all viewpoints ${\bf x}_{i}$ along a path $\mathcal{P}_{j}$:
$\operatorname*{arg\,min}_{\mathcal{P}_{j}}\sum_{{\bf x}_{i}\in\mathcal{P}_{j}}C_{\textit{local}}(\text{Log}({\bf x}_{i}))-\eta\log(\mathtt{EIG}_{\mathcal{P}_{j},i}({\bf x}_{i}))$ (13)
where $\eta$ is a hyper-parameter controlling the importance between EIG and localization accuracy. The agent can then explore the environment with planned path $\mathcal{P}$. Our active mapping system constantly updates the map, and we replan using our active path planning algorithm if we detect the agent is getting close to a possible obstacle or upon reaching the end of the previously selected path.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
ANS 16.34 0.6818 0.3923 0.3886 0.1105 10.49
Active-INR (gt) 22.66 0.7652 0.2164 0.1528 - 9.20
UPEN (gt) 21.31 0.7325 0.2714 0.1696 - 8.79
UPEN 16.44 0.6678 0.4134 0.4841 0.5158 8.58
ExplORB 18.99 0.7175 0.3994 0.2664 0.2296 9.00
FBE 21.45 0.7618 0.2126 0.1028 0.1680 10.64
Ours 23.28 0.8067 0.2507 0.0696 0.0226 10.88
Table 1: Quantitative Evaluation on Reconstruction Quality and Tracking Accuracy on Gibson dataset
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
UPEN (gt) 15.58 0.5175 0.3936 0.4548 - 12.69
UPEN 12.23 0.4795 0.5157 0.7356 0.4393 10.72
FBE 15.80 0.5952 0.4392 0.4085 1.2004 15.69
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Table 2: Quantitative Evaluation on Reconstruction Quality and Tracking Accuracy on HM3D dataset
Refer to caption Refer to caption
Ground truth Estimated
Figure 6: Qualitative Comparison for Trajectories on Cantwell from Gibson Dataset. Left is the ground truth trajectory, right is the estimated trajectory.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
w.o. LLM & Localization Uncertainty 16.15 0.6550 0.6193 0.3409 0.2478 16.06
w.o. LLM 16.94 0.6799 0.5847 0.2887 0.1694 17.26
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Table 3: Ablation Study of Localization Uncertainty Term on Scenes from the HM3D Dataset. We compare our method with and without the localization uncertainty term to validate that including it provides improvements on both localization and reconstruction
4 Experiments
4.1 Experimental Set-up
Dataset
Following previous methods [22, 85], our algorithm is evaluated in the Habitat Simulator [73] on the Gibson [81] and Habitat-Matterport 3D (HM3D) [56] datasets comprised of indoor scenes reconstructed from scans of real houses. We use all the scenes in the val split for the Gibson dataset. We also evaluate our method on single-floor scenes whose scene ID can be found in the supplementary. We adopt the default start point in the Habitat Simulator as the starting point for exploration in each scene. The total number of steps for each experiment is 2000. The system takes color and depth images at the resolution of 800x800 and outputs a discrete action at each step. The action space consists of MOVE FORWARD by 5cm, TURN LEFT, and TURN RIGHT by 5°$\degree$. The field of view (FOV) is set to 90°$\degree$ vertically and horizontally. Please refer to the supplement for more details about the evaluation split and other hyper-parameters.
Metrics
We evaluated our method using the Peak-signal-to-noise ratio (PSNR), Structural Similarity Index Measure  (SSIM) [79], Learned Perceptual Image Patch Similarity (LPIPS) [89] for RGB rendering and mean absolute error (MAE) for depth rendering as metrics for scene reconstruction quality. We calculate these metrics using 2000 points uniformly sampled from the movement plane of the agent in the scene, discarding any points that are not navigable. We argue that the rendering quality reflects both reconstruction quality and pose accuracy because high tracking accuracy would help the training of the 3D Gaussian Splatting model. Meanwhile, misaligned poses will lead to misaligned rendering at test time thus leading to inferior results. Following previous approaches [22, 7], we also use coverage in $m^{2}$ as an evaluation metric. To evaluate the pose estimation accuracy, we use the root mean squared average tracking error (RMSE ATE), but as the trajectories for each method are different, the RMSE ATE should only be considered along with other metrics such as coverage.
Baselines
We compare to two exploration methods which assume ground truth pose: UPEN [22] and Active Neural Mapping (active-INR) [85]. UPEN chooses trajectories to reduce the epistemic uncertainty of the occupancy prediction produced by an ensemble of models. Active-INR aims to minimize the prediction robustness against random weight perturbation of its signed distance field scene representation. We also compare our method with Active Neural SLAM (ANS) [7], explORB [53] and Frontier Based Exploration (FBE) [82] without ground truth pose provided. ANS [7] learns to predict a map and estimate a global goal based on it. ExplORB [53] computes the Fisher Information of the Hessian on its pose graph optimization.
To compare the rendering quality fairly, we run all the baselines using the MonoGS [44] backend for reconstruction. We run UPEN and FBE online, but for ANS, active-INR, and ExplORB, we record and playback trajectories obtained using their source code. Because the forward step size for ANS is much larger than for our method, we interpolate the trajectory so that the forward step size matches our method’s to make the steps comparable. For ExplORB, since the official implementation is based on MoveBase, which uses velocity commands, we sample the trajectory at 5 Hz. We also found that ANS, active-INR, and UPEN failed on some scenes due to localization failure. ANS produces a pose estimate (using information from noisy pose sensors not provided to our pipeline), so we set the pose estimate of the MonoGS backend to the one from ANS. As active-INR and UPEN do not produce a pose estimate, we evaluate them using the ground-truth pose.
4.2 Comparison Against Previous Methods
Table 1 shows the results of our method and the baselines for exploration in scenes from the Gibson dataset [81], and Table 2 shows the results of our method and some baselines on HM3D [6]. Our method outperforms the baselines on all metrics. We further qualitatively compare the reconstruction qualities after active exploration in Fig. 4 and Fig. 5, and the trajectories in Fig. 6. Our method does not have major errors from failed localization, and we have fewer gaps in the scenes than other methods. For example, in the Ribera scene, all methods except for us and FBE miss the bathroom at the bottom left, and FBE misses more of the area around the sofa than us. For the trajectories, we show the estimated and ground truth trajectories for the Cantwell scene from the Gibson dataset. Cantwell is a relatively large and challenging scene, so it is suitable for showing the differences between methods. We show only a few baselines to keep the figure legible. We can see that Active-INR stays in a smaller area than the rest of the methods. ANS often goes close to walls, whereas FBE and our method are generally more toward the center of corridors or rooms. ANS also extends further into rooms than these methods. FBE does not go to the region on the bottom right, and ANS does not go to the region on the bottom left, whereas our method efficiently visits most areas of the scene.
4.3 Ablation Study
To validate the effectiveness of localization uncertainty and LLM, we performed an ablative study of each component in our system on the Gibson dataset in Table. 3. As can be seen, the average trajectory error is much lower with the localization uncertainty than without. In addition, after adding LLM, the model performs better in the overall rendering. The ATE slightly increases after adding LLM, but this could be attributed to the larger area traversed by the robot.
5 Conclusion
We present an active mapping system that can autonomously explore an environment, the first method to do so using a 3D Gaussian representation without ground truth camera poses. By rendering maps using our scene representation, we can seamlessly query a multimodal Large Language Model for long-horizon planning to leverage its knowledge of scene layouts. This allows the robot to efficiently explore the scene while avoiding setting unreachable navigation goals. We then plan an optimal path to the long-horizon goal using our information-driven path proposal and selection algorithm, which balances the information gain with respect to the map against the localization uncertainty, ensuring that the robot captures the geometric details of the scene for a high-quality reconstruction. We evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D  [56] datasets in terms of the rendering quality, coverage, and average tracking error. We compare to recent exploration methods with [22, 85] and without [53, 7] ground-truth pose, using the same scene representation for reconstruction as us for fairness, and show that our method has superior performance. To enable our method to support more robotics applications, future work could extend it to take advantage of our 3D scene representation and consider movement with higher degrees of freedom (DOF) than the currently supported 3DOF. Incorporating semantic features [91, 68, 55] to allow for grounding language to the scene would also enable many robotics and computer vision applications.
Acknowledgements
Financial support by the grants NSF FRR 2220868, and ONR N00014-22-1-2677 is gratefully acknowledged.
Appendix A Appendix
In this supplementary material, we discuss things we left over in our main paper due to page constraints. We provide more background information about 3DGS and the derivative with respect to localization parameters in Sec. A.1. We discussed more implementation details in Sec. A.2 and provided the prompt we use for the Multimodal LLM and an example of the interaction in Sec. A.5 to help reproduce our results. The source code of this project will be made public soon. We also include detailed quantitative and qualitative results in Sec. A.3 and Sec. A.4.
A.1 Additional Background Information
For the completeness of our method, we also provide the key definition for the 3D Gaussian Splatting backbone [33] and 3D Gaussian SLAM [44]. In 3D Gaussian Splatting [33], the rendered pixel color is calculated by composing all 3D Gaussians projected in a tile.
$\hat{C}(\text{r})=\sum_{i=1}^{N_{s}}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right){\bf c}_{i}$ (14)
$\alpha_{i}=\text{exp}\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right)(1-\text{exp}(-\sigma_{i}\delta_{i}))$ (15)
$\delta_{i}=t_{i+1}-t_{i}$ represents the distance between adjacent samples, and $N_{s}$ indicates the number of samples. ${\bf c}_{i}$ is the color of each 3D Gaussian given the current view direction ${\bf d}$ and $\sigma_{i}$ is given by evaluating a 2D Gaussian with covariance $\Sigma$.
The Jacobian of the localization parameters are defined as:
$\frac{\mathcal{D}{\boldsymbol{\mu}_{C}}}{\mathcal{D}{{\bf x}}}=\begin{bmatrix}\boldsymbol{I}&-\boldsymbol{\mu}_{C}^{\times}\end{bmatrix}\text{ and }\frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{{\bf x}}}=\begin{bmatrix}\mathbf{0}&-\mathbf{W}_{:,1}^{\times}\\ \mathbf{0}&-\mathbf{W}_{:,2}^{\times}\\ \mathbf{0}&-\mathbf{W}_{:,3}^{\times}\\ \end{bmatrix}~{},$ (16)
where × denotes the skew symmetric matrix of a 3D vector, and $\mathbf{W}_{:,i}$ refers to the $i$th column of the matrix.
Unlike PSNR, the EIG can be computed without ground truth images, making it possible to perform view selection during exploration.
A.2 Implementation Details
The 2D occupancy map’s resolution is 5cm. For each single frontier pixel on the 2D map, we add 200 3D Gaussians, which are uniformly distributed in the 3D cube above it. Other parameters like color, opacity, and scale are generated uniformly between 0 and 1. When there are frontiers on the 2D map, we choose the next frontier to be explored by the area of each frontier divided by the distance. When no frontier exists, we select the top 20% of Gaussians with the highest score. These Gaussians are grouped using DBSCAN [19]. The largest cluster is selected for candidate pose generation. Candidates are uniformly sampled in the range between 0.3m to 1m, facing towards the selected position. Only the poses in free space are kept for path-level selection. The importance factor $\eta$ in Eq. 13 is set to 5 across all experiments. The source code for this project will be made public no later than the publication of this paper.
We compute the Expected Information Gain (EIG) for each global candidate and use A* to plan a path for each of them. In order to prevent a twisted path, we consider locations 0.15m (3 pixels) away from the current robot position as neighbors and set the robot width to 3 pixels for collision check. However, the path planned by A* might have redundant waypoints, causing unnecessary turns for the robot. Therefore, we smooth the path by finding shortcuts. Specifically, for each waypoint $w_{i}$, if the path between waypoint $w_{i+2}$ and $w_{i}$ is collision-free, then we remove the intermediate waypoint $w_{i+1}$ from the path. Finally, we use a greedy follower for motion planning. If the angle between the heading direction of the robot and the relative next waypoint is larger than 5°$\degree$, then we turn left or right to decrease the angle. Otherwise, we choose the forward action to approach the next waypoint. In such a way, we get a sequence of actions $\{a_{i}\}_{i=1}^{T}$ for each path.
Given a sequence of actions $\{a_{i}\}_{i=1}^{T}$ for each path, we use forward dynamics to compute the future camera poses $\{c_{i}\}_{i=1}^{T}$. Initially, we use an intermediate variable ${\mathbf{H}}^{\prime\prime}_{\text{obs}}\triangleq{\mathbf{H}}^{\prime\prime}[{\bf w}^{*}]$ to help compute expected information gain along the path. For each camera pose $x_{i}$, we compute its pose Hessian ${\mathbf{H}}^{\prime\prime}_{\text{pose}}$ and the current model Hessian matrix ${\mathbf{H}}^{\prime\prime}_{\text{cur}}\triangleq{\mathbf{H}}^{\prime\prime}[{\bf y}|x_{i},{\bf w}^{*}]$. ${\mathbf{H}}^{\prime\prime}_{\text{cur}}$ is then accumulated, and we update ${\mathbf{H}}^{\prime\prime}_{\text{obs}}$ to evaluate the remaining poses on the path. We select the path that minimizes the objective given by Eq. 13 for execution.
The compute of Fisher Information in our experiments takes 30 frames per second on our Nvidia L40 GPU. The time on path planning will not significantly impact the robot’s motion if the path is planned ahead of time and runs in parallel as the robot moves. The regularizer $\lambda$ in Eq. 4 is $10^{-6}$ across all experiments. The weight parameters $\eta$ in Eq. 13 is 0.2 across all experiments.
A.3 Results for Each Scene in Gibson and HM3D Dataset
Following previous literature [85], we use the following scenes for Gibson Dataset: Greigsville, Denmark, Cantwell, Eudora, Pablo, Ribera, Swormville, Eastville, Elmira. For HM3D we use the following scenes: DBjEcHFg4oq, mscxX4KEBcB, QKGMrurUVbk, oPj9qMxrDEa, CETmJJqkhcK. The detailed results for each scene on each evaluation metric are presented as bar plots in Fig. 7 for Gibson and Fig. 8 for HM3D. We also present qualitative comparisons on testing views from the Gibson dataset in Fig. 9.
A.4 Detailed version of qualitative results
We provide larger versions of the qualitative rendering comparisons from the main paper, Fig. 11 shows the Gibson scenes and Fig. 12 shows the HM3D scenes.
A.5 Example of Using Multimodal LLM
We provide an example of our interaction with the multimodal LLM in Fig 13.
1
2<system> You are an AI assistant that can analyze images and plan a long-term goal for the exploration task of a ground robot.
3You will be given a bird-eye view image of a scene.
4The goal is to plan a long-term exploration mission for a robot to traverse the area.
5The robot's task is to explore the terrain efficiently, identifying important areas, potential obstacles, and unvisited areas.
6Please analyze the image and select a long-term goal from the candidates for the robot to explore the area.
7Empty space doesn't always mean they are unvisited regions, sometimes it's just outside the floor plan of this scene.
8We are allowed to explore a total of <TOTAL\_STEPS> steps and this is step <STEP\_ID>.
9Therefore, it's better to select a space that is close to the visited regions but still unvisited and not behind the walls.
10The current location of the robot is marked with the blue star(*) marker.
11The last frontier you selected is marked with a yellow diamond shape.
12The visited path is painted as green lines in the image.
13Note that you don't have to select the closest point to the robot, but the point that is most likely to be unvisited and important to explore.
14As you can see, there are <NUM\_FRAME> candidate points to select from.
15They are numbered from 0 to <NUM\_FRAME - 1> in red color.
16If you find all the goals are not necessary to explore and we should instead focus on improving existing reconstruction, please give -1 in the `target` entry of the JSON.
17Please provide a detailed exploration plan and select an exploration target with reasons in the JSON format as shown below.
18
19
20```
21{
22"target": 2, "reason": "The target is located at an unvisited region of the image and seems to be an unvisited bedroom"
23}
24```
25Do not cut off the JSON and generate the full JSON.
26</system>
27
28
29<user>: I have a bird-eye view image of a scene. The goal is to plan a long-term exploration mission for a robot to traverse the area. Please analyze the attached image and provide the exploration plan first and then an exploration target in the specified JSON format.
30</user>
Figure 7: Per-scene results on Gibson Dataset
Figure 8: Per-scene results on HM3D Dataset
ANS Refer to caption Refer to caption Refer to caption
Active-INR (gt) Refer to caption Refer to caption Refer to caption
UPEN Refer to caption Refer to caption Refer to caption
ExplORB Refer to caption Refer to caption Refer to caption
FBE Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 9: Test Rendering Qualitative Visualization on Gibson Dataset All the renderings are from the test view of the Gibson dataset.
UPEN Refer to caption Refer to caption Refer to caption
FBE Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption
Figure 10: Test Rendering Qualitative Visualization on Habitat-Matterport 3D Dataset All the renderings are from the test view of the Habitat-Matterport 3D dataset.
ANS [7] Active-INR [85] UPEN [22]
Refer to caption Refer to caption Refer to caption
ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
ANS [7] Active-INR [85] UPEN [22]
Refer to caption Refer to caption Refer to caption
ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative Comparison for Final Scene Reconstruction on Gibson Dataset Greigsville (top) and Ribera (bottom) scenes. We provide top-down rendering for different methods. Note that UPEN and Active-INR use GT pose in this visualization.
UPEN [22] FBE [82] Ours
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: Qualitative Comparison for Final Scene Reconstruction on Habitat-Matterport 3D Dataset mscxX4KEBcB (top), oPj9qMxrDEa (middle) and QKGMrurUVbk (bottom) scenes. We provide top-down rendering for different methods.
Figure 13: An Example about a Dialog with the Multi-modal LLM. The chat box on the top is the question from our method, and the bottom is the response from GPT-4o, which we parsed to get a long-term goal for our path proposal. We provided the contextual information in both the textual prompt and the image.